OpenStudy (anonymous):

Past experience shows that a process requires $17.50 worth of material for every $12 spent on labour. How much should be budgeted for material if the budget for labour is $17,760?

OpenStudy (xapproachesinfinity):

USE PROPORTIONS \[$12=$17 ~~Worth ~of ~material\] \[17760=x\] so \[\frac{x}{17}=\frac{17760}{12} \Longrightarrow x=\frac{17760}{12}17\]

OpenStudy (xapproachesinfinity):

use ur calculator :)

OpenStudy (anonymous):

thanks :)